# Transformers for Stratified Spectropolarimetric Inversion: Proof of Concept

### v1

@@ 3.4 Training and validation
For both the transformer and MLP models, we generate a total of 1 million unique training examples from a single snapshot of the MANCHA simulation, consisting of synthetic Stokes profiles and their corresponding atmospheric models. These data are randomly split into training and validation sets ($85\%/15\%$). The generalisation and test performance are evaluated on a completely independent snapshot from the same simulation, taken 90 seconds later. Both the transformer and MLP are trained and evaluated on identical datasets to enable a direct comparison.
@@
Hyperparameter searches are performed for both architectures: for the transformer, we vary learning rate (0.001, $1\times 10^{-4}$, $3\times 10^{-4}$), hidden dimension (128, 256, 384), number of encoder and decoder layers (2, 3, 4), and number of attention heads (2, 4). For the MLP baseline, we explore learning rates (0.001, $1\times 10^{-4}$, $3\times 10^{-4}$), hidden dimensions (128, 256, 384, 512), and the number of layers (3, 4, 6, 8). The batch size is primarily determined by available GPU memory, with typical values of 32–64. The best model for each architecture is selected as that with the lowest validation loss during training. For the MLP, the lowest validation loss was achieved with a learning rate of $3\times 10^{-4}$, a hidden dimension of 512, and 8 layers. For the transformer, the lowest validation loss was achieved with a learning rate of $3\times 10^{-4}$, a hidden dimension of 128, 4 layers, and 4 attention heads.
@@ 4.1 Loss curves
Figure 5: Training (solid lines) and validation (dashed lines) loss curves for the transformer (blue) and MLP (red) models, as a function of epoch. The vertical blue dotted line indicates the epoch at which the transformer achieves its lowest validation loss (epoch 34, validation loss 0.136), while the vertical red dotted line shows the early stopping point for the MLP (epoch 32, validation loss 0.1749). Both models show smooth convergence, with no evidence of overfitting.
Before benchmarking inversion performance, we first examine the training and validation loss curves for both the transformer and MLP models, shown in Figure 5. For each architecture, we plot the MSE on the training and validation sets as a function of epoch for the best-performing hyperparameter configuration. In both cases, the loss curves show smooth convergence, with validation loss plateauing or increasing after a certain epoch, indicating effective early stopping and an absence of overfitting or memorisation. This ensures that subsequent comparisons are made using well-generalised models.
@@ 4.2 Generalisation testing and comparing the transformer and MLP predictions
Figure 6 shows maps of the temperature at selected optical depths, comparing the ground truth, transformer prediction, MLP prediction, and the residuals. At $\log(\tau)=0$, the transformer accurately reproduces the surface granulation pattern, capturing even subtle small-scale features within intergranular lanes. The MLP performs similarly well across most of the field of view, but exhibits pronounced errors in several regions—particularly within intergranular lanes—where it severely underestimates the temperature and fails to converge on a physically plausible solution. The residual maps confirm that the transformer outperforms the MLP throughout, with the largest MLP errors exceeding 200 K, whereas transformer errors in the same regions remain typically below 200 K, reflecting greater stability. These trends persist at $\log(\tau)=-1$, with the transformer consistently delivering more accurate and robust predictions. However, at $\log(\tau)=-4$, both models show increased absolute residuals and systematic errors, though the transformer still captures the overall temperature morphology. The MLP retains its tendency for localised, severe mispredictions but otherwise exhibits comparable stability to the transformer at this optical depth.
@@
For both magnetic field strength (Fig. 7) and line-of-sight velocity (Fig. 8), the qualitative trends mirror those seen in the temperature maps. At lower optical depths ($\log(\tau)=0$ and $-1$), the transformer accurately reconstructs the spatial morphology and small-scale features across the field of view, with residuals typically lower than those of the MLP. The transformer especially predicts a solution closer to ground truth for the weakest magnetic fields. The MLP performs well in much of the domain, but is prone to conspicuous localised failures, leading to large under- or over-estimation in those areas. The residual maps highlight the superior stability and consistency of the transformer, with the MLP occasionally exhibiting extreme errors. At $\log(\tau)=-4$, both models show larger and more systematic deviations, but the transformer continues to better reproduce the overall morphology, while the MLP is again affected by scattered zones of severe misprediction. For magnetic field strengths at higher layers both models noticeably struggle more than in the photosphere. The transformer more successfully predicts the strongest magnetic field strengths at the higher layer, but also fails to predict the largest downflows.
@@
For inclination (Fig. 9), the transformer is a clear leader in terms of reproducing the highly complex small-scale morpgological features at all three optical depths. The MLP continues to suffer from very poor convergence in the same pixels as the other parameters.
@@
The comparative performance of the two models is further illustrated in Figure 10, which presents 2D histograms of the predicted versus ground truth values at two representative optical depths for all pixels in the test snapshot, with the density of points indicated by a logarithmic colour scale. This means there are order(s) of magnitude more pixels along the brightest regions of the histograms. The dashed red line corresponds to perfect one-to-one agreement. For all parameters at both depths, the transformer predictions cluster more tightly along the one-to-one line, indicating not only high accuracy but also a strong absence of systematic bias. In contrast, the MLP results show greater scatter in all parameters and both depths. For instance, the scatter of temperature values at $\log(\tau)=-4$ shows a systemtic error for the MLP that the transformer does not suffer from. For inclination in the photosphere, the transformer has a much higher density of correct predictions along the one-to-one correspondence line, while the MLP is more broadly dispersed—this pattern is also evident for temperature and line-of-sight velocity. This confirms that the transformer model delivers more reliable and stable inversions across atmospheric heights, however both models show the greatest differences emerging in the chromospheric regime.
Correlations are not useful for examining the true error between the predictions of the transformer and MLP and the ground truth, but they are useful as a measure of performance because if one model is outperforming the other its correlations should be significantly higher. Figure 11 shows the correlation coefficients as a function of optical depth for the transformer and the MLP for each parameter. At first we computed the Pearson correlation coefficient, $r$, but found that the correlations were typically and surprisingly near-zero for the MLP. The $r$ is extremely sensitive to outliers, so we additionally calculated the Spearman rank coefficient, $\rho$. From the $\rho$ one can observe that both the transformer and MLP achieve similar highly positively correlated values ($\rho$ above 0.8) in the photosphere for temperature and line-of-sight velocity. For temperature, the $\rho$ for the MLP drops between $\log(\tau)=-2$ and $\log(\tau)=-3$ while the transformer remains stable and highly correlated. The same behaviour occurs for line-of-sight velocity. At even higher layers, the $\rho$ significantly drops (below 0.8) for both models in both temperature and velocity between $\log(\tau)=-4$ and $\log(\tau)=-5$.
@@
It is clear that for both models and the magnetic parameters the inference above $\log(\tau)=-5$ is unreliable. However, the $\rho$ values for temperature and line-of-sight velocity are observed to degrade above $\log(\tau)=-5$ and rise again, especially for line-of-sight velocity, between $\log(\tau)=-6$ and $\log(\tau)=-7$.
@@ 4.3 Impact of Gaussian noise on inversion performance
Figure 12 shows the $r$ as a function of optical depth for the transformer when trained at different noise levels, ranging from $5\times 10^{-4}$ Ic to $5\times 10^{-2}$ Ic. Since the transformer does not produce extreme failed predictions like the MLP we do not need to calculate the $\rho$ in order to examine the impact of noise. The noise degradation has no significant impact on the inference of temperature between $log(\tau)=1$ and $log(\tau)=-4$. In the most extreme noise cases, some minor degradation in the $r$ value begins at around $log(\tau)=-4$ before rapidly dropping at $log(\tau)=-5$. The lower noise cases ($1\times 10^{-4}$ Ic, $1\times 10^{-3}$ Ic) retain comparably high $r$ values relative to the noiseless case. The inference of line-of-sight velocities does begin to degrade above $log(\tau)=-2$ for the higher noise values, but the two lowest noise levels never significantly fall below the noiseless benchmark and indeed even out-perform the noiseless case at $log(\tau)=-5$ for the lowest noise levels.
@@ 4.4 Transformers as a regulariser
Figure 13: Comparison of ground truth atmospheric parameters transformer and MLP predictions for a pixel near an intergranular-granule boundary. The upper row shows a close-up view of $T$, $B$, $v_{\mathrm{LOS}}$, and $\gamma$ ground truth values at $log(\tau)=-1$. Shown in the left column is the Stokes vector produced by DeSIRe from the ground truth atmospheric model, for the selected pixel, as a function of spectral pixel index. Shown in the right column is the atmospheric parameters as a function of optical depth. The black curve shows the ground truth, the dashed blue curve shows the transformer prediction, and the red curve shows the MLP prediction.
To qualitatively demonstrate the performance of both ML models, we selected a pixel near the intergranular-granule boundary and compared the predicted atmospheric parameters and the corresponding synthetic Stokes profiles. Figure 13 presents the ground truth atmospheric model parameters alongside the MLP and transformer predictions as a function of optical depth, as well as the Stokes I, Q, U, and V profiles produced from the ground truth model - the MLP and transformer have predicted the model parameters from this Stokes vector. The Stokes vector contains very weak chromospheric polarisation signals, and the MLP and transformer both correctly predict very low $B$ values in the chromosphere. From the Stokes vector, one can observe some Stokes $V$ profiles in the photospheric lines have opposite signs or appear nearly single lobed compared to others. This is indicative of a line-of-sight velocity or inclination gradient, which is indeed present in the ground truth. For line-of-sight velocity, the ground truth has a clear gradient in optical depth which both the MLP and transformer correctly predicts. For inclination, the transformer prediction very closely matches the ground truth in the photosphere, but the MLP produces an unphysical, oscillatory solution, which violates plausible values above $180^{\circ}$.
@@
Figure 14 shows inference for a pixel in a relatively strong magnetic concentration located in an intergranular lane. The Stokes vector for this model has larger circular polarisation amplitudes in the upper photospheric and chromospheric lines. As in the previous case, the transformer and MLP both track the ground truth temperature closely in the photosphere and up to $log(\tau)=-4$, however above this both ML predictions deviate significantly from the true values. The same is true for line-of-sight velocity, although the transformer has a lower residual error even in the photosphere. For magnetic field strength the transformer clearly predicts more accurately than the MLP, as the MLP hallucinates variations in the photosphere that are not present in the ground truth. Consistently with the previous example, the MLP struggles significantly when predicting inclination, even predicting unphysical values (below $0^{\circ}$ in this case) while the transformer prediction stays physically bounded.
We emphasise that these two examples are not those which clearly shows an unphysical predicted solution from the MLP maps (Fig. 6 to  9). In those pixels, the model parameters predicted by the MLP are orders of magnitude away from the ground truth and are objectively physically implausible.
@@ 5 Discussion
In this proof of concept study, we have demonstrated the application of a transformer-based neural network, originally developed for natural language processing, to the regression problem of inferring stratified atmospheric parameters from full-Stokes, multi-line spectropolarimetric data generated from realistic 3D MHD simulations. Our results show that the transformer model outperforms an MLP baseline in nearly all aspects: reproducing the spatial and vertical morphology of temperature, magnetic field strength, and velocity; achieving higher correlations with the ground truth at all optical depths; and, crucially, delivering physically plausible solutions even in challenging chromospheric layers where traditional methods and the simpler ML model cannot be trusted.
@@
One of the most significant findings is the stability of the transformer’s predictions in both the photosphere and towards the chromosphere. While the MLP baseline performs at least comparably to the transformer in the deeper atmospheric layers in temperature and velocity, it is nevertheless prone to severe localised failures, sometimes producing unphysical or extreme outputs. In contrast, the transformer is markedly more robust, accurately capturing the fine structure of both granules and intergranular lanes and providing sensible solutions even where the input polarisation signal is weak. Positional embedding allows the transformer to remain regularised in its predictions even in these challenging regimes.
@@
The attention mechanism at the core of the transformer architecture enables the model to leverage non-local spectral features, potentially corresponding to Zeeman splitting or subtle asymmetries, which are often overlooked by models that flatten or ignore the spectral dimension. Although a detailed interpretability study is beyond the current scope, initial inspection of the attention maps suggests that the model prioritises physically meaningful regions of the spectrum (especially the most magnetically sensitive Fe I lines) for inference. This could open new avenues for understanding how neural networks extract physical parameters from complex solar spectra.
@@
Despite these promising results, several limitations and caveats should be acknowledged explicitly. First, the models are trained and evaluated exclusively on synthetic data produced by the MANCHA simulation code. While this simulation is one of the most realistic options for quiet Sun conditions due to its inclusion of the Biermann battery effect, its chromosphere is not expected to be realistic. The “ground truth” in this context is therefore only as realistic as the underlying simulation. Extension to simulations such as Bifrost, which might provide a more realistic chromosphere, or to training sets that combine multiple simulation codes, will be necessary to ensure that the model can generalise to a broader set of solar conditions. However, expanding the training set to include models and Stokes vectors from multiple simulation codes with differing physical assumptions risks introducing degenerate mappings between Stokes profiles and atmospheric parameters. This can hinder model convergence and reduce generalisation, unless carefully controlled. In particular, differences in chromospheric realism and magnetisation between simulations may lead to conflicting training signals that cannot be reconciled by the model.

### v2

@@ 3.4 Training and validation
For both the transformer and MLP models, we generate a total of 1 million unique training examples from a single snapshot of the MANCHA simulation, consisting of synthetic Stokes profiles and their corresponding atmospheric models. These data are randomly split into training and validation sets ($85\%/15\%$). The generalisation and test performance are evaluated on a completely independent snapshot from the same simulation, taken $450$ seconds later. Both the transformer and MLP are trained and evaluated on identical datasets to enable a direct comparison.
@@
Hyperparameter searches are performed for both architectures: for the transformer, we vary learning rate (0.001, $1\times 10^{-4}$, $3\times 10^{-4}$), hidden dimension (128, 256, 384), number of encoder and decoder layers (2, 3, 4), and number of attention heads (2, 4). For the MLP baseline, we explore learning rates (0.001, $1\times 10^{-4}$, $3\times 10^{-4}$), hidden dimensions (128, 256, 384, 512, 1024, 2048), and the number of layers (3, 4, 6, 8). The batch size is primarily determined by available GPU memory, with typical values of 32–64. The best model for each architecture is selected as that with the lowest validation loss during training. For the MLP, the lowest validation loss was achieved with a learning rate of $3\times 10^{-4}$, a hidden dimension of 2048, and 8 layers. For the transformer, the lowest validation loss was achieved with a learning rate of $3\times 10^{-4}$, a hidden dimension of 128, 4 layers, and 4 attention heads.
We also considered the impact of dropout on the performance of the MLP. The transformer applies a dropout rate of $0.1$ within its feed-forward layers and residual connections, as per the default PyTorch implementation. We added dropout layers after each activation in the MLP architecture, using a dropout rate of $0.1$ to match the default setting in the Transformer’s encoder and decoder blocks. The validation loss increased significantly, suggesting that dropout was not beneficial for the MLP in this setting.
@@
The code used in this project to train the machine learning models is open source and is available444www.github.com/r-j-campbell/SINN-inversions [^5] https://doi.org/10.5281/zenodo.16793363.
@@ 4.1 Loss curves
Figure 5: Training (solid lines) and validation (dashed lines) loss curves for the transformer (blue) and MLP (red) models, as a function of epoch. The epoch at which the transformer achieves its lowest validation loss is epoch 34 with a validation loss of 0.112. The lowest validation loss for the MLP was at epoch 34 with 0.1417.
Before benchmarking inversion performance, we first examine the training and validation loss curves for both the transformer and MLP models, shown in Figure 5. For each architecture, we plot the MSE on the training and validation sets as a function of epoch for the best-performing hyperparameter configuration. In both cases, the loss curves show smooth convergence, with validation loss plateauing, indicating effective early stopping and an absence of overfitting. This ensures that subsequent comparisons are made using well-generalised models.
@@ 4.2 Generalisation testing and comparing the transformer and MLP predictions
Figure 6 shows maps of the temperature at selected optical depths, comparing the ground truth, transformer prediction, MLP prediction, and the residuals. At $\log(\tau)=0$, the transformer accurately reproduces the surface granulation pattern, capturing even subtle small-scale features within intergranular lanes. The MLP performs similarly well across most of the field of view, but exhibits larger errors than the transformer, particularly within intergranular lanes. The residual maps confirm that the transformer outperforms the MLP throughout, with the largest MLP errors exceeding 200 K, whereas transformer errors in the same regions remain typically below 200 K, reflecting greater stability. These trends persist at $\log(\tau)=-1$, with the transformer consistently delivering more accurate and robust predictions. However, at $\log(\tau)=-4$, both models show increased absolute residuals and systematic errors, though the transformer still better captures the overall temperature morphology.
@@
For both magnetic field strength (Fig. 7) and line-of-sight velocity (Fig. 8), the qualitative trends mirror those seen in the temperature maps. At lower optical depths ($\log(\tau)=0$ and 1-1), the transformer accurately reconstructs the spatial morphology and small-scale features across the field of view, with residuals typically lower than those of the MLP. The transformer especially predicts a solution closer to ground truth for the weakest magnetic fields. The residual maps highlight the superior stability and consistency of the transformer. At $\log(\tau)=-4$, both models show larger and more systematic deviations, but the transformer continues to better reproduce the overall morphology. For magnetic field strengths at higher layers both models noticeably struggle more than in the photosphere. The transformer more successfully predicts the strongest magnetic field strengths at the higher layer, but also fails to predict the largest up-flows.
@@
For inclination (Fig. 9), the transformer is a clear leader in terms of reproducing the highly complex small-scale morphological features at all three optical depths.
@@
The comparative performance of the two models is further illustrated in Figure 10, which presents 2D histograms of the predicted versus ground truth values at two representative optical depths for all pixels in the test snapshot, with the density of points indicated by a logarithmic colour scale. This means there are order(s) of magnitude more pixels along the brightest regions of the histograms. The dashed red line corresponds to perfect one-to-one agreement. For all parameters at both depths, the transformer predictions cluster more tightly along the one-to-one line, indicating not only high accuracy but also a strong absence of systematic bias. In contrast, the MLP results show greater scatter in all parameters and both depths. For inclination in the photosphere, the transformer has a much higher density of correct predictions along the one-to-one correspondence line, while the MLP is more broadly dispersed. This pattern is also evident for temperature and line-of-sight velocity. This confirms that the transformer model delivers more reliable and stable inversions across atmospheric heights, however both models show the greatest differences emerging towards the chromospheric regime.
Correlations are not useful for examining the true error between the predictions of the transformer and MLP and the ground truth, but they are useful as a measure of performance because if one model is outperforming the other its correlations should be significantly higher. Figure 11 shows the Pearson correlation coefficients, $r$, as a function of optical depth for the transformer and the MLP for each parameter. One can observe that both the transformer and MLP achieve similar highly positively correlated values ($r$ close to $1$) in the photosphere for temperature and line-of-sight velocity. For temperature, the $r$ drops for the MLP above $\log(\tau)=-3$, while the transformer remains stable and highly correlated. At even higher layers, the $r$ remains above 0.8 for the transformer in temperature between $\log(\tau)=-3$ and $\log(\tau)=-6$.
@@
The $r$ for inclination, azimuthal parameters, and to a lesser extent the line-of-sight velocity, shows a dip between $\log(\tau)=-2$ and $\log(\tau)=-3$. This may represent a gap in sensitivity between the chromospheric Ca II lines and the Fe I photospheric lines. One may have expected the upper photospheric lines to fill this gap, but the very weak polarisation signals in these lines is a limiting factor.
@@ 4.3 Impact of Gaussian noise on inversion performance
Figure 12 shows the $r$ as a function of optical depth for the transformer when trained at different noise levels, ranging from $5\times 10^{-4}$ Ic to $5\times 10^{-2}$ Ic. The noise degradation has no significant impact on the inference of temperature between $\log(\tau)=1$ and $\log(\tau)=-4$. In the most extreme noise cases, some minor degradation in the $r$ value begins at around $\log(\tau)=-4$, and all noise levels rapidly deteriorate at $\log(\tau)=-6$. The inference of line-of-sight velocities only begins to degrade above $\log(\tau)=-4$ for the higher noise values, but the two lowest noise levels never significantly fall below the noiseless benchmark.
@@ 4.4 Transformers as a regulariser
Figure 13: Comparison of ground truth atmospheric parameters transformer and MLP predictions for a pixel near an intergranular-granule boundary. The upper row shows a close-up view of $T$, $B$, $v_{\mathrm{LOS}}$, and $\gamma$ ground truth values at $\log(\tau)=-1$. Shown in the left column is the Stokes vector produced by DeSIRe from the ground truth atmospheric model, for the selected pixel, as a function of spectral pixel index. Shown in the right column is the atmospheric parameters as a function of optical depth. The black curve shows the ground truth, the dashed blue curve shows the transformer prediction, and the red curve shows the MLP prediction.
To qualitatively demonstrate the performance of both ML models, we selected a pixel near the intergranular-granule boundary and compared the predicted atmospheric parameters and the corresponding synthetic Stokes profiles. Figure 13 presents the ground truth atmospheric model parameters alongside the MLP and transformer predictions as a function of optical depth, as well as the Stokes $I$, $Q$, $U$, and $V$ profiles produced from the ground truth model - the MLP and transformer have predicted the model parameters from this Stokes vector. The Stokes vector contains very weak chromospheric polarisation signals, and the MLP and transformer both correctly predict very low $B$ values in the chromosphere. For temperature, both models make excellent predictions. For line-of-sight velocity, the ground truth has a clear gradient in optical depth which both models correctly predict. For the two magnetic parameters shown, inclination and magnetic field strength, the transformer prediction matches the stratification of the ground truth in the photosphere much more closely.
@@
Figure 14 shows inference for a pixel in a relatively strong magnetic concentration located in an intergranular lane. The Stokes vector for this model has larger circular polarisation amplitudes even in the upper photospheric and lower chromospheric lines. As in the previous case, the transformer and MLP both track the ground truth temperature closely in the photosphere, however above this both ML predictions deviate significantly from the true values. The same is true for line-of-sight velocity, although the transformer has a lower residual error even in the photosphere. For magnetic field strength the transformer clearly predicts more accurately than the MLP. Consistently with the previous example, the MLP struggles significantly when predicting inclination, even predicting unphysical values (both above $180^{\circ}$ in this case) while the transformer prediction stays physically bounded, even though it has a significant error above the photosphere.
@@ 5 Discussion
In this proof of concept study, we have demonstrated the application of a transformer-based neural network, originally developed for natural language processing, to the inversion problem of inferring stratified atmospheric parameters from full-Stokes, multi-line spectropolarimetric data generated from realistic 3D MHD simulations. Our results show that the transformer model outperforms an MLP baseline in nearly all aspects: reproducing the spatial and vertical morphology of temperature, magnetic field strength, and velocity; achieving higher correlations with the ground truth at all optical depths; and, crucially, delivering physically plausible solutions even in challenging chromospheric layers.
@@
One of the most significant findings is the improvement of the transformer’s predictions in the magnetic parameters. While the MLP baseline performs at least comparably to the transformer in the deeper atmospheric layers in temperature and velocity, the transformer is markedly more robust at accurately capturing the fine structure of both granules and intergranular lanes, clearly outperforming the MLP in the inference of magnetic parameters, even where the polarisation signal is weak. Positional embedding allows the transformer to remain regularised in its predictions even in these challenging regimes, while the MLP violates physically plausible bounds in its predictions in inclination, and we have verified this is not an isolated incident for the MLP.
@@
The attention mechanism at the core of the transformer architecture enables the model to leverage non-local spectral features, potentially corresponding to Zeeman splitting or subtle asymmetries, which may be overlooked by models that treat each wavelength as just another input feature, with no explicit bias toward local or long-range spectral relationships. Although a detailed interpretability study is beyond the current scope, initial inspection of the attention maps suggests that the model prioritises physically meaningful regions of the spectrum (especially the most magnetically sensitive Fe I lines) for inference. This could open new avenues for understanding how neural networks extract physical parameters from complex solar spectra.
@@
Despite these promising results, several limitations and caveats should be acknowledged explicitly. First, the models are trained and evaluated exclusively on synthetic data produced by the MANCHA simulation code. While this simulation is one of the most realistic options for quiet Sun conditions due to its inclusion of the Biermann battery effect, its chromosphere is not expected to be realistic. The “ground truth” in this context is therefore only as realistic as the underlying simulation. Extension to simulations which might provide a more realistic chromosphere, or to training sets that combine multiple simulation codes, will be necessary to ensure that the model can generalise to a broader set of solar conditions. However, expanding the training set to include models and Stokes vectors from multiple simulation codes with differing physical assumptions risks introducing degenerate mappings between Stokes profiles and atmospheric parameters. This can hinder model convergence and reduce generalisation, unless carefully controlled. In particular, differences in chromospheric realism and magnetisation between simulations may lead to conflicting training signals that cannot be reconciled by the model.